OpenStudy (orange_):

Question:

OpenStudy (orange_):

Five friends want to rent a go-cart track that costs $100. They already have $20. How much more does each friend need to contribute if they each contribute the same amount? Let x represent any unknown value in the problem. Select each correct answer. When writing an equation to solve this problem, x should represent the amount each friend needs to contribute. When writing an equation to solve this problem, x should represent the total amount the friends need to contribute. When writing an equation to solve this problem, x should represent the total cost to rent the go-cart track. The equation 5x+20=100 can be used to solve this problem. The equation 5(x+20)=100 can be used to solve this problem. The equation x+20=100 can be used to solve this problem.Each friend needs to contribute $80. Each friend needs to contribute $16. Each friend needs to contribute $4. Each friend needs to contribute $80.

OpenStudy (orange_):

FAN AND MEDAL!!

OpenStudy (orange_):

Please help me!

OpenStudy (orange_):

@Awolflover1

OpenStudy (orange_):

@YoungStudier @TheSmartOne

OpenStudy (orange_):

What?

OpenStudy (orange_):

Oh, nvm

OpenStudy (timamits):

From the step 5x + 20 = 100 to 5(x+20) =100 and then x+20 =100 are wrong steps. you need to proceed from 5x + 20 = 100 to 5x = 100-20=80 and then x = 80/5.

OpenStudy (timamits):

Just tell how much total amount they need to contribute?

OpenStudy (orange_):

I just need to find which ways would equal how much they need to contribute

TheSmartOne (thesmartone):

So they need a total of 100 dollars. They already have 20. The friends need to pitch in equal amount of money to make up for the remaining. Let's say each friend pitches in x dollars. 5x is the remaining amount of money they need to pitch in, besides the 20 dollars, to make 100 dollars. Can you write an equation for that?